5. A porpoise jumps straight up and crashes back into the water at 8.9 m/s. The drag and buoyancy forces of the water slow the porpoise down with an acceleration of -9.3 m/s2 as the porpoise finally slows to a stop. Find the depth the porpoise reaches.
1
Expert's answer
2020-09-23T09:00:29-0400

The distance traveled by a porpoise yo the

"d=\frac{v^2-v_0^2}{2a}=\rm\frac{0^2-8.9^2}{2\times (-9.3)}=8.5\: m"